The Mann Company belongs to a risk class for which the appropriate discount rate is 12 percent. The company currently has 222Please Show all work and formulas
Answer #1

First understand what is the significance of discount rate of 12%.
This also means that expected return on the stock is 12%.

If no dividend is declared price of stock after 1 year will be = Price Today * (1 + expected return)

So price of stock in case of no dividend = 114*(1 + 12%) = 127.68

Now, if dividend to be paid at the end of the year is $4, share price will reduce by $4.
In case of dividend share price = 127.68 - 4 = 123.68

Alternatively, Using standard formulas -
P0 = (P1 + D1)/(1 + R) to get the same answer.


For dividend case -
144 = (P1 + 4)/1.12
P1 = 123.68

For case of no dividend -
144 = (P1 + 0)/1.12
P1 = 127.68


Part 1:

$123.68

Part 2:

$127.68

Part 3:

Immediate outflow on investments = $4.7 million
Net Income for the year = $2.1 million
Dividend to be paid = $4 per share
Number of shares outstanding = 222,000
Current Share Price = $114
Appropriate Discount Rate = 12%

As company needs the funding at the beginning of the year, when share price is $114
Number of shares to be issues = $4.7 million/$114 = 41,228.07

Hence company should issue 41,229 shares.